# Multinucleon transfer with time-dependent covariant density functional theory

### 12 Jan 2024 v1

@@ Acknowledgements.
We thank S. Szilner for useful discussions. This work has been supported in part by the High-end Foreign Experts Plan of China, the National Natural Science Foundation of China (Grants No. 11935003, 11975031, 12070131001, and 12141501), the High-performance Computing Platform of Peking University, the QuantiXLie Centre of Excellence, a project co-financed by the Croatian Government and European Union through the European Regional Development Fund - the Competitiveness and Cohesion Operational Programme (KK.01.1.1.01.0004), and the Croatian Science Foundation under the project Uncertainty quantification within the nuclear energy density framework (IP-2018-01-5987). We acknowledge the funding support from the State Key Laboratory of Nuclear Physics and Technology, Peking University.

### 22 Feb 2024 v2

@@ Acknowledgements.
We thank S. Szilner for useful discussions. This work has been supported in part by the High-end Foreign Experts Plan of China, the National Natural Science Foundation of China (Grants No. 11935003, 11975031, 12070131001, and 12141501), the High-performance Computing Platform of Peking University, the project “Implementation of cutting-edge research and its application as part of the Scientific Center of Excellence for Quantum and Complex Systems, and Representations of Lie Algebras”, PK.1.1.02, European Union, European Regional Development Fund, and by the Croatian Science Foundation under the project Relativistic Nuclear Many-Body Theory in the Multimessenger Observation Era (IP-2022-10-7773). We acknowledge the funding support from the State Key Laboratory of Nuclear Physics and Technology, Peking University.